OpenStudy (anonymous):

For the function rule y = -2x - 3, provide the x-y coordinates of the y-intercept and the x-intercept. would i just graph this then the two meets that's where the y-intercept and x-intercept is?

OpenStudy (anonymous):

if you want the x-intercept put y=0 and solve for x if you want the y-intercept put x=0 and solve for y

OpenStudy (anonymous):

okay, can you check my answer. for the first one 0=-2x-3 x=-3/2 ( kinda of a weird answer) 2nd one y=-2*0-3 y=-3

OpenStudy (anonymous):

you got it

OpenStudy (anonymous):

x-intercept (-3/2;0) y-intercept (0;-3)

OpenStudy (anonymous):

got it. Thanks a bunch!!

OpenStudy (anonymous):

yw